OpenStudy (anonymous):

Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)? g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x − 10 − 5x g(x) − f(x) = −2x − 8 − 3x g(x) − f(x) = −2x − 10 − 3x @ganeshie8 plz help

OpenStudy (texaschic101):

g(x) - f(x) 3x - 9 - (5x - 1) = 3x - 9 - 5x + 1 = -2x - 8

OpenStudy (texaschic101):

I am not the best with functions...you might better get ganshie in here

OpenStudy (anonymous):

i think it's g(x)-f(x)=3x-8-5x @iamabarbiegirl